OpenStudy (anonymous):

integrate x^3dx/(x^2-a^2)^3 .... thanks

OpenStudy (anonymous):

Geez this is a long one....

OpenStudy (anonymous):

r u done with this ....or should i give try?

OpenStudy (anonymous):

I posted the solution already

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

could u try?

OpenStudy (anonymous):

@heromiles: is that the answer already??

OpenStudy (anonymous):

yeah y don't u add and subtract a^3 and then try solvin it

OpenStudy (anonymous):

lol thats a copy from wolframa

OpenStudy (anonymous):

also you can just sub u^2 =x^2 -a^2 so 2u du = 2x dx

OpenStudy (anonymous):

u du = x dx so = [(u^2+a^2) /(u^6) ] u du

OpenStudy (anonymous):

= \[= \int\limits \frac{u^2+a^2}{u^5} = \int\limits( u^{-3} + a^2 u^{-5} )du \]

OpenStudy (anonymous):

\[= -\frac{1}{2u^2} - \frac{a^2}{4u^4} +C = - \frac{1}{ 2(x^2-a^2)} - \frac{a^2}{4(x^2-a^2)^2} +C\]

OpenStudy (anonymous):

which is equivalent to the answer wolframa gives, but its 10times faster, didnt involve cheating and is alot more elegant :)